OpenStudy (anonymous):

Suppose a parabola has vertex (-8,-7) and also passes through the point (-7,-4). What is the equation of the parabola in vertex form

OpenStudy (jack1):

just spit-balling here, but you could try simultaneous equations for this: y = ax^2 + bx + c -4 = 49a -7b + c ENQ 1 also -7 = 64a - 8b +c EQN 2 and y' = 2ax + b 0 = -16a + b EQN 3

OpenStudy (jack1):

... so b = 16a subbed into EQN 2 -7 = 64a - 8b +c -7 = 64a - 8(16a) +c -7 = 64a - 128a + c -7 = -64a + c so c = -7 + 64a

OpenStudy (anonymous):

@Jack1 you used information for y' , where did you get that?

OpenStudy (jack1):

subbed into Equation 1 -4 = 49a -7b + c ENQ 1 -4 = 49a -7b -7 + 64a -4 = 49a -7(16a) -7 + 64a -4 = 49a -112a -7 + 64a -4 = a -7 so a = 3 so 0 = -16a + b 0 = -48 + b b = 48

OpenStudy (jack1):

c = -7 + 64a c = -7 + 64*3 c = -7 + 192 c = 185

OpenStudy (jack1):

@mathessentials y' = derivative of the line at the vertex, y' = turning point so gradient = 0

OpenStudy (anonymous):

I see :) it makes sense, thanks for explaining

OpenStudy (jack1):

all good dude

OpenStudy (jack1):

anyways, i think i went the wrong way, this gives you the equation of the line in the form \[y = ax^2 + bx + c\] in this case - \[y = 3x^2 + 48x + 185\] whereas I think vertex form is something like: \[y = a(x – h)^2 + k\] where "k" is ... the y value of the vertex? and "h" = the x value? I know "a" is the same in either... so your vertex form equation should look something more like \[y = 3(x + 8)^2 -7\] ... maybe...?